OpenStudy (anonymous):

Let R be the region enclosed by the curve \[y = 1 / (r^2 + 1)\] and r-axis from 0 to 1. Find the volume of the solid generated when the Region R is revolved about the r-axis

OpenStudy (anonymous):

what's r axis?

OpenStudy (anonymous):

its the question :)

OpenStudy (anonymous):

whats the answers i also want to know too

OpenStudy (anonymous):

i'm working on it :) i don'T have the answer yet too

OpenStudy (anonymous):

\[V = \int\limits_{0}^{1}\pi ( 1 / (r^2 + 1))^2dx \]

OpenStudy (anonymous):

ok so the formula is \[V=\pi \int\limits_{0}^{1}f(r)^2 dr\] so \[V=\pi \int\limits_{0}^{1}1/(r^2+1)^2 dr\] using the substitution r=tan(theta) you find \[dr=\sec ^2(\theta) d \theta\] so \[V=\pi \int\limits\limits_{0}^{\pi/4}\sec^2\theta/(\tan ^2\theta+1)^2 d \theta\] which is \[\pi(2+\pi)/8\]

OpenStudy (anonymous):

yeap i got the same answer too :)